OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

Parth (parthkohli):

Determine the slope of (3,7) and (-2,-8). A parallel line will have the same slope.

OpenStudy (anonymous):

How do I determine slope? What is the equation I would use?

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {y_2 - y_1 \over x_2 - x_1} }\)

OpenStudy (binary3i):

if two lines are parelel they have thew same slope.

OpenStudy (anonymous):

So the slope is three?

Parth (parthkohli):

y2 is -8 y1 is 7. x2 is -2 x1 is 3.

Parth (parthkohli):

1/3 actually

Parth (parthkohli):

So then the same slope would be the parallel line's slope. So 1/3

OpenStudy (anonymous):

Ohh! I put y^1 - y^2 and x^1 - x^2 by mistake. And thank you so much, this is seriously a lifesaver right now !

OpenStudy (binary3i):

now y=mx+c where you have m=3 y=3x+c yous the point to get c 3=-2*3 +c c=9 so the line is y=3x +9

OpenStudy (anonymous):

Thanks!

Parth (parthkohli):

Actually it's 3 oops

OpenStudy (anonymous):

Woah! I got it, okay, Messed up negative signs. It's 3!

Parth (parthkohli):

I got confused. It's 3....GRR!

OpenStudy (anonymous):

Thanks! (: